# The Unreasonable Ineffectiveness of the Deeper Layers

### v1

@@ 4.1 Accuracy on QA benchmarks
Our first set of results are shown in Figure 2, where we plot 555-shot MMLU accuracy as a function of the fraction of layers removed: in the left panel we present the Llama-2 family, in the middle panel we present models from the Qwen family, and in the right panel we show Mistral-7B and Phi-2. In order to better compare models of different total number of layers, in these plots we opted to normalize the $x$-axis by the fraction of layers removed (rather than the absolute number of layers removed). Note that since MMLU contains multiple choice questions with four possible responses, the expected accuracy of random guessing is 25%.
@@
Importantly, we see a characteristic flat region of robust performance followed by a sharp transition to random accuracy at a pruning fraction around 45%-55% for models in the Llama-2 family, 35% for Mistral 7B, 25% for Phi-2, and 20% for models from the Qwen family. This implies that the essential knowledge required to achieve a model’s top score isn’t removed by significant layer removal – even though the fraction can be quite large(!) – until eventually that knowledge is lost at a critical model-dependent threshold. [^6] This effect is rather robust to choice of QA benchmark: in Appendix Figure 6 we plot the average 0-shot BoolQ accuracy for our model families and observe analogous behavior. Contrasting the curves with and without healing, we see that finetuning offers a modest improvement by better preserving the unpruned performance and pushing the phase transition to random guessing to slightly larger pruning fractions.
@@ 4.2 Loss on next-token predictions
In this section, we look at the effect of layer pruning on the pretraining optimization objective – the cross-entropy loss of next-token prediction – when evaluated on a subset of the C4 validation dataset. [^7] We make sure that none of the validation data are seen during the healing stage. In order to have a fair comparison across models with different sized vocabularies $V$, we normalize the loss by $\log V$, which corresponds to the loss of sampling tokens randomly with uniform probability. (See Appendix A.2 for more details.)
In Figure 3 , we plot the normalized C4 validation loss for all seven of our models, after healing (left panel) and before healing (right panel), as a function of the fraction layers removed. Without healing, we see that there is a somewhat sharp(ish) transition to random guessing for each model at approximately the pruning fraction that the QA benchmark accuracies also sharply transition to random guessing, suggesting that models are hopelessly harmed at this point, cf. Figure 2. Next, contrasting the scales of both plots, we see that healing significantly restores the next-token prediction ability of all the models to near-unpruned levels, with the loss increasing slowly and linearly with layer dropping. Most strikingly – from a scientific perspective – is the post-healing continuity through the pruning fractions where we previously found sharp transitions for the QA benchmarks: this decoupling illustrates one way of disconnecting (or creating a miscalibration) between performance on downstream tasks – such as MMLU and BoolQ – and continuous measures of performance – such as the cross-entropy loss.  [^8] This is consistent with Ref. Schaeffer et al. (2023) that argued jumps in one kind of metric may not be visible in others.
@@ 4.3 Angular distances between representations
Recall from earlier Figure 1(c): for Llama-2-70B this plotted the angular distance $d(x^{(\ell)},x^{(\ell+n)})$ that compared the $\ell$-th layer to the $(\ell+n)$-th layer, across all initial indexes $\ell$ for block sizes from $n=1$ to $n=64$; the minimum of the curves, $\ell^{\star}(n)$, gave the optimal block to prune for a given $n$, cf. (6). A more compact way to display this same data is shown in the heat maps of Figure 4: each square is colored to depict the row-normalized angular distance between layer $\ell$ and $\ell+n$ across all possible $\ell$, and $n$ up to very large fractions of the total number of layers; the optimal layer to prune for a given block size, $\ell^{*}(n)$, corresponds to the minimal distance in each row.
@@
Figure 4: Normalized angular distance (7) from initial layer $\ell$ (x-axis) with block size $n$ (y-axis) for each of the seven models we evaluated; the distance for each $n$ is shifted and rescaled to span the same range, [0,1][0,1][ 0 , 1 ] (yellow to purple): the optimal block to prune, $\ell^{*}(n)$, corresponds to the deepest yellow for each row. Across models, the deeper layers tend to be very similar, though the deepest blocks that include the final layer (squares along the outer diagonal) are (near-)maximally dissimilar.
@@ 4.4 A simpler pruning strategy
[…]
@@ 5 Discussion and Future Directions
Beginning with the release of the open-weight LLaMA family (Touvron et al., 2023b), the open-source machine-learning community has rallied around the philosophy of making LLMs accessible to everyone. This has engendered many innovations around efficiency, such as LoRA Hu et al. (2021) and quantization (with LoRA) Dettmers et al. (2023), allowing large (near-)state-of-the-art 70B models to be finetuned on only single 80GB A100 GPUs. In conjunction with these other tools, our work enables further efficiency gains via a simple-to-implement layer-pruning technique.
In particular, the released version of Llama-2-70B spans 140140140 GB of memory and consumes approximately $3\times 10^{10}$ FLOPs per token. With 4-bit quantization and a layer-pruning fraction of 50%, the model fits in approximately 17.517.517.5 GB of memory and requires roughly $1.5\times 10^{10}$ FLOPs per token: quantization from 16-bit bfloats to 4-bit QLoRA precision reduces model memory by a factor of 4, but keeps FLOPs more or less the same, since calculations are performed in 16-bit precision; layer pruning will additionally reduce both memory and FLOPs by an amount equal to the layer-pruning fraction.These memory and compute requirements enable open-weight state-of-the-art models to be run and even finetuned efficiently on consumer-level GPUs without any CPU off-loading and with only minor performance trade-offs.
At the conclusion of the work, we are left with the following questions:
@@
Some of these questions would benefit from studying both layer similarity and pruning across different pretraining checkpoints; for instance, at what point does the sharp phase transition and critical depth in the QA accuracies emerge, and does more training lead to better use of the prunable parameters? Others suggest explorations with different pretraining architectures and objectives, e.g. in order better make use of the deeper layers. With more comprehensive evaluations, if different kinds of tasks degrade at very different depths, then this might indicate that the knowledge required to complete those tasks is stored at different depths. [^10] Alternatively, one could measure $d(x^{(\ell)},x^{(\ell+n)})$ or find $\ell^{*}(n)$ as a function of different eval datasets. It would be very interesting to use pruning to systematically study these kind of interpretability questions.
@@
Appendix A Experimental Details
Here we explain various details of models and healing (§A.1) and of evaluations (§A.2).
A.1 Model and healing details
@@
  • For quantization, we used the bitsandbytes library for QLoRA Dettmers et al. (2023) to quantize our models to 4 bits.
@@
  • Also following Ref. (Lee et al., 2023), we only applied LoRA to FFN modules: ["gate_proj", "down_proj", "up_proj"] for Llama-2 and Mistral models, ["fc1", "fc2"] for Phi-2, and ["w1", "w2", "c_proj"] for Qwen models.
The large majority of these hyperparameter choices are standard and found in previous works, e.g. Refs. (Lee et al., 2023; Dettmers et al., 2022). For absolute clarity, we list display all the model specific architecture and healing details below:

### v2

@@ 4.1 Accuracy on QA benchmarks
Our first set of results are shown in Figure 2, where we plot $5$-shot MMLU accuracy as a function of the fraction of layers removed: in the left panel we present the Llama-2 family, in the middle panel we present models from the Qwen family, and in the right panel we show Mistral-7B and Phi-2. In order to better compare models of different total number of layers, in these plots we opted to normalize the $x$-axis by the fraction of layers removed (rather than the absolute number of layers removed). Note that since MMLU contains multiple choice questions with four possible responses, the expected accuracy of random guessing is 25%.
@@
Importantly, we see a characteristic flat region of robust performance followed by a sharp transition to random accuracy at a pruning fraction around 45%-55% for models in the Llama-2 family, 35% for Mistral 7B, 25% for Phi-2, and 20% for models from the Qwen family. This implies that the essential knowledge required to achieve a model’s top score isn’t removed by significant layer removal – even though the fraction can be quite large(!) – until eventually that knowledge is lost at a critical model-dependent threshold. [^4] This effect is rather robust to choice of QA benchmark: in Figure 7 we plot the average 0-shot BoolQ accuracy for our model families and observe analogous behavior. Contrasting the curves with and without healing, we see that finetuning offers a modest improvement by better preserving the unpruned performance and pushing the phase transition to random guessing to slightly larger pruning fractions.
@@ 4.2 Loss on next-token predictions
In this section, we look at the effect of layer pruning on the pretraining optimization objective – the cross-entropy loss of next-token prediction – when evaluated on a subset of the C4 validation dataset. [^5] We make sure that none of the validation data are seen during the healing stage. In order to have a fair comparison across models with different sized vocabularies $V$, we normalize the loss by $\log V$, which corresponds to the loss of sampling tokens randomly with uniform probability. (See Appendix B.2 for more details.)
In Figure 3 , we plot the normalized C4 validation loss for all seven of our models, after healing (left panel) and before healing (right panel), as a function of the fraction layers removed. Without healing, we see that there is a somewhat sharp(ish) transition to random guessing for each model at approximately the pruning fraction that the QA benchmark accuracies also sharply transition to random guessing, suggesting that models are hopelessly harmed at this point, cf. Figure 2. Next, contrasting the scales of both plots, we see that healing significantly restores the next-token prediction ability of all the models to near-unpruned levels, with the loss increasing slowly and linearly with layer dropping. Most strikingly – from a scientific perspective – is the post-healing continuity through the pruning fractions where we previously found sharp transitions for the QA benchmarks: this decoupling illustrates one way of disconnecting (or creating a miscalibration) between performance on downstream tasks – such as MMLU and BoolQ – and continuous measures of performance – such as the cross-entropy loss.  [^6] This is consistent with Schaeffer et al. (2023) that argued jumps in one kind of metric may not be visible in others.
@@ 4.3 Angular distances between representations
Recall from earlier Figure 1(c): for Llama-2-70B this plotted the angular distance $d(x^{(\ell)},x^{(\ell+n)})$ that compared the $\ell$-th layer to the $(\ell+n)$-th layer, across all initial indexes $\ell$ for block sizes from $n=1$ to $n=64$; the minimum of the curves, $\ell^{\star}(n)$, gave the optimal block to prune for a given $n$, cf. (6).
A more compact way to display this same data is shown in the heat maps of Figure 4: each square is colored to depict the row-normalized angular distance between layer $\ell$ and $\ell+n$ across all possible $\ell$, and $n$ up to very large fractions of the total number of layers; the optimal layer to prune for a given block size, $\ell^{*}(n)$, corresponds to the minimal distance in each row.
@@
Figure 4: Normalized angular distance (7) from initial layer $\ell$ (x-axis) with block size $n$ (y-axis) for each of the seven models we evaluated; the distance for each $n$ is shifted and rescaled to span the same range, $[0,1]$ (yellow to purple): the optimal block to prune, $\ell^{*}(n)$, corresponds to the deepest yellow for each row. Across models, the deeper layers tend to be very similar, though the deepest blocks that include the final layer (squares along the outer diagonal) are (near-)maximally dissimilar.
@@ 4.4 A simpler pruning strategy
Inspired by our recent conclusions, we experiment with a very simple heuristic pruning strategy: (1) if pruning $n$ layers from an $L$-layer model, drop layers $(L-n)$ to $(L-1)$ so as to remove the deepest block that excludes the final layer; then (2) heal with a small amount of finetuning as before. Compared with our principal similarity-informed pruning strategy, this simpler heuristic algorithm has the advantage of never requiring practitioners to load onto a GPU or inference the unpruned model. It also provides a meaningful ablation of the importance of optimizing the block to prune.
@@ 5 Discussion and Future Directions
At the end of this work, many readers are puzzled by the following: are the deeper layers entirely useless? So far, we’ve provided evidence that the elimination of the deeper layers does not affect performance on QA tasks like MMLU (Figure 2), while at the same time have shown that their removal does disrupt the next-token predictions of the underlying model (Figure 3). Since perplexity often correlates with performance on downstream tasks, which are the tasks that are hurt by layer pruning?
Here are two hypotheses consistent with the fact that the model’s perplexity is disturbed proportionally to pruning fraction:
  • (i)
    The deeper layers are not essential for storing knowledge, but are useful for more complicated computations, such as those that involve reasoning.
  • (ii)
    The deeper layers are necessary when the model has to generate many tokens before answering a question, such as when it produces a chain-of-thought (CoT).
We test these hypotheses by evaluating our layer-pruned models on tasks that involve CoTs or reasoning. For the former, we’ll look at Chain-of-Thought MMLU (CoT-MMLU); for the latter, we’ll look at GSM8K (Cobbe et al., 2021), a grade-school math benchmark, and HellaSwag (Zellers et al., 2019), a multiple choice common-sense reasoning benchmark. [^7] Here are the details for how we performed these three evaluations: For CoT-MMLU, we followed the flan_cot_fewshot evaluation in EleutherAI (Gao et al., 2023), in which models produce a chain of thought before generating their answer. Note that the accuracy at $0\%$ pruning fraction for MMLU without CoT is much better than the analogous accuracy at $0\%$ pruning fraction for CoT-MMLU ($\sim 69\%$ vs. $\sim 43\%$, respectively; cf. Figures 2 and 6), consistent with some previous work (e.g., see Table 16 of Chung et al. (2024)). For GSM8K, we used the gsm8k_cot evaluation in EleutherAI (Gao et al., 2023) and measured pass@1; for each problem we extracted an answer from a single generation (with CoT) and checked for correctness against the ground-truth answer. For HellaSwag, we used the hellaswag evaluation in EleutherAI (Gao et al., 2023). Note that HellaSwag is a multiple-choice benchmark, so random performance is 25%.
In Figure 6, we plot the performance of Llama-2 70B pruned with the similarity-informed pruning strategy across CoT-MMLU (left), GSM8K (center), and HellaSwag (right): on the one hand, both GSM8K and HellaSwag, our two reasoning tasks, exhibit immediate degradation in performance with any amount of pruning, correlating with a similar decrease in the perplexity evals (Figure 3); on the other hand, CoT-MMLU shows a relatively flat region of robust performance with pruning, analogous to our previous results on QA benchmarks (e.g. Figure 2). This is some initial evidence for hypothesis (i) over hypothesis (ii): the deeper layers may be useful for higher-level reasoning tasks, while less important for knowledge intensive QA tasks; moreover, perplexity errors due to pruning do not compound to hurt QA evals when the model is required to generate many tokens.
Figure 6: Evaluation of Llama-2 70B with the similarity-informed pruning strategy across different evaluation tasks. (Left: Chain-of-Thought MMLU (CoT-MMLU), Center: GSM8K, Right: HellaSwag.) We see that GSM8K and HellaSwag show immediate degradation of performance with any level of pruning, while CoT-MMLU behaves qualitatively similarly to MMLU without CoT; this suggests that the deeper layers are likely necessary for reasoning tasks.
Now at the conclusion of the work, we are left with the following questions:
@@
  • Can we devise a pruning strategy that is robust for reasoning tasks?
@@
Some of these questions would benefit from studying both layer similarity and pruning across different pretraining checkpoints; for instance, at what point does the sharp phase transition and critical depth in the QA accuracies emerge, and does more training lead to better use of the prunable parameters? Others suggest explorations with different pretraining architectures and objectives, e.g. in order better make use of the deeper layers (for example, one can imagine applying layer dropout (Fan et al., 2019) or early exit during pre-training (Elhoushi et al., 2024) to induce equal usage of layers). With more comprehensive evaluations, if different kinds of QA tasks degrade at very different depths, then this might indicate that the knowledge required to complete those tasks is stored across different layers. [^9] Alternatively, one could measure $d(x^{(\ell)},x^{(\ell+n)})$ or find $\ell^{*}(n)$ as a function of different eval datasets. It would be very interesting to use pruning to systematically study these kind of interpretability questions.
@@
Appendix A Extended Literature Review
In this section, we review practical strategies for post-training efficiency and discuss some scientific investigations that provide motivation for, or insight into, our approach: in §A.1, we first review the history of pruning and then discuss its modern application to LLMs; in §A.2, we contrast pruning with distillation, an alternative strategy for reducing the parameter count of LLMs; then in §A.3, we discuss the various practical methods for efficient finetuning and inference acceleration that can be used in conjunction with our pruning strategy; finally in §A.4 we highlight some scientific investigations into some depth-dependent statistical properties of LLMs that are complementary to our results.
A.1 Pruning
Pruning is a method for reducing the size of a trained machine-learning model by removing unnecessary parameters, either individually or together as a group. Pruning for neural networks has a long history (LeCun et al., 1989, Hassibi and Stork, 1992), and, as originally conceived, unstructured pruning techniques sparsify networks by removing individual parameters based on pre-defined criteria. For instance, if a parameter of the model has a very small value, then removing it – i.e. by setting it to exactly zero – will likely have minimal impact on performance. Inspired by this early work, modern researchers began exploring different criteria for such unstructured pruning, focusing mostly on computer vision models (Han et al., 2015, Chen et al., 2015, Srinivas and Babu, 2015). In particular, Han et al. (2015) developed an iterative pruning method for alternatively pruning and finetuning a network in order to reach better compression ratios and performance.
While these models were smaller, they were not necessarily more efficient: sparsifying networks by removing individual parameters according to a criterion leads to irregular or pseudorandom sparsification patterns that are difficult to accelerate without specialized hardware or libraries designed for sparsity (Li et al., 2016). To that end, structured pruning techniques were developed to remove irrelevant groups of parameters together, such as particular channels or filters in convolutional networks. As this increased their practical relevance, researchers then began exploring structured pruning across computer vision (Li et al., 2016, Wen et al., 2016, Hu et al., 2016, He et al., 2017, Huang et al., 2018) and pre-transformer NLP architectures (Murray and Chiang, 2015, See et al., 2016, Kim and Rush, 2016).
Following unprecedented progress in language modeling, recent work has focused on applying structured pruning methods to the Transformer (Vaswani et al., 2017). These studies consider nearly every possible component of the model architecture for elimination, with methods ranging from dropping attention heads (Voita et al., 2019, Michel et al., 2019, Kim and Awadalla, 2020), to dropping layers (Fan et al., 2019, Zhang and He, 2020, Fan et al., 2021, Jha et al., 2023, Sajjad et al., 2023, Liu et al., 2023a), to pruning hidden states (Hou et al., 2020), to rank reducing large weight matrices (Sharma et al., 2023), replacing sparse weight matrices with smaller dense ones (Ashkboos et al., 2024), to many combinations of the aforementioned groups (Xia et al., 2022, Lagunas et al., 2021).
Of the prior work that also considers transformer layer dropping, most (Fan et al., 2019, Zhang and He, 2020, Fan et al., 2021, Xia et al., 2022, Sajjad et al., 2023) study BERT-style models (Devlin et al., 2018), while we consider decoder-only GPT-style models (Radford et al., 2019) that are most commonly used for large-scale language modeling and generation. BERT-style models are naturally suited for understanding tasks due to their bidirectional masked language modeling (MLM) objective, while GPT-style models are instead suited for generation, due to their autoregressive objective. While this divide has been questioned in light of more powerful GPT-style models (Zhong et al., 2023), previous work (Ethayarajh, 2019) has found significant qualitative differences between BERT and GPT models in terms of the evolution of the layer-wise representation of words. Altogether, this suggests that layer-dropping strategies will behave differently between the two families.
One study for BERT-style pre-trained models, Sajjad et al. (2023), concludes that the best layer-pruning strategy is dropping the final layers; this partially resonates with our results, although in contrast we find that (a) for some pruning sizes keeping the last few layers of the model is actually beneficial, and that (b) for all pruning sizes keeping the very last layer is essential. Additionally, while the authors also study similarity between representations in different layers – as in our approach – they actually found a higher similarity between representations in the shallow layers compared to the deeper ones – which very sharply disagrees with our results. Importantly, the models considered in Sajjad et al. (2023) consist of a few hundred million parameters, which is much smaller than the model scales we consider in our work. Perhaps as a consequence, the authors didn’t observe the sharp transition in downstream accuracies that we report in §4.1, despite the fact that they also finetuned their pruned models.
In contrast, while Jha et al. (2023) does consider GPT-style models, the methodology is quite different from ours: (i) rather than pretraining first and then using a fixed layer-dropping strategy as we do, instead the authors incrementally drop layers in a modified pretraining procedure; and (ii) the authors study their own sub-1B parameter models, while we focus on the families of readily available, open-weight, large-scale 2.7B-70B parameter models that are commonly used and/or finetuned for practical applications.
As we were finalizing our preprint, Men et al. (2024) was posted: this paper empirically studies different layer-pruning strategies for GPT-style models (Llama-2 7B and Baichuan2-7B-base) and their subsequent effects on benchmarks (MMLU, CMMLU, and CMNLI). They investigate various layer-importance metrics – notably, their "Block Influence" function is similar to our cosine similarity metric – and find that they are able to prune up to $\sim$28of layers of Llama-2 7B with minimal impact on performance. This provides independent evidence supporting our main takeaway that the deeper layers are not critical for storing knowledge.
Finally, a systematic approach to layer dropping in transformers has also been studied in the context of wav2vec models, which are encoder-only models that map speech to embeddings and are sized in the hundred-million parameter regime (Baevski et al., 2020). With these models, Liu et al. (2023a) developed a layer-pruning algorithm based on the correlation between layers and downstream metrics. Beyond the model architecture and domain, one significant difference between this and our work is that Liu et al. (2023a) considered non-contiguous pruning proposals, e.g. dropping alternate layers. Our intuition for layer pruning predicts that this shouldn’t work as well – at least for decoder-only language models – as it creates multiple mismatches, one with each block of layers removed.
A.2 Model distillation
A completely different method for reducing the size of a trained machine-learning model is model distillation (Hinton et al., 2015), in which knowledge is transferred from a large “teacher” model to a smaller “student” model by training the student on the distribution predicted by the teacher. The essential insight is that this can transform the very general knowledge and capabilities of the teacher into more streamlined, compressed, and possibly skill-specific representations.
While a very general technique, in the setting of language models, distillation has been implemented with (a) white-box approaches, in which the the student is trained to imitate the teacher’s logits (Gu et al., 2023) or hidden states (Jiao et al., 2019); as well as with (b) black-box approaches, in which the student only has access to the output tokens generated by the teacher. This latter approach broadly covers cases where the student is trained on text that is augmented by the teacher in some way, such as by adding synthetic labels (Wang et al., 2021), generating high quality synthetic text (Eldan and Li, 2023, Li et al., 2023a, Gunasekar et al., 2023) by providing chain of thought reasoning (Fu et al., 2023, Hsieh et al., 2023), which aims to enhance the student’s reasoning skills, or by annotating instructions that enhance the student’s instruction-following capabilities (Jiang et al., 2023b).
A.3 Efficient finetuning and inference acceleration
Complementary to directly reducing size of a model, parameter-efficient finetuning (PEFT) focuses on reducing the cost of specializing LLMs to certain tasks. In particular, Low Rank Adapters (LoRA) reduce the memory and compute of fine tuning by freezing the pretrained model and introducing a parametrically small number of additional trainable weights (Hu et al., 2021). We use its quantized cousin, QLoRA (Dettmers et al., 2023), to keep our experiments cost efficient. Other PEFT methods that can be combined with our work are  Li et al. (2023b) and Zhang et al. (2023): in the first, the initialization of the LoRA matrices is adjusted to a quantization scheme; in the second, LoRA ranks for different LLM modules are chosen in an adaptive manner.
For additional efficiency gains we could combine our layer-pruned models with methods that further accelerate inference: with speculative decoding (Leviathan et al., 2023), tokens are rapidly generated from a smaller draft model and then evaluated in parallel by the main model; with Medusa (Cai et al., 2024) the draft model is discarded for extra decoding heads, but ultimately achieves a similar effect. In particular, it could be interesting to consider highly-compressed layer-pruned models as potential draft models in a speculative decoding setup.
A.4 A breadth of depth-dependent studies
Finally, let us highlight some scientific work that study the depth-dependent properties of LLMs. One relevant direction considers how knowledge and linguistic properties are encoded in language models. On the one hand, Meng et al. (2022) and Dai et al. (2021) analyze the storage and recall of factual associations: these works emphasize that knowledge localizes within the middle (Meng et al., 2022) or final (Dai et al., 2021) layers, which has implications for directly editing or erasing part of a model’s factual knowledge. On the other hand, attempts to perform such editing gives evidence that information may be stored non-locally across layers (Hase et al., 2023). Relatedly, Geva et al. (2023) investigates the way facts are processed during inference, distinguishing between the role of attention heads, for attribute extraction, and the MLP blocks, for subject enrichment: both are delocalized across several layers.
Next, following the earlier “logic lens” (nostalgebraist, 2020), Belrose et al. (2023) invented a technique they called “tuned lens” to study the trajectory of predictions by using a learnable affine transformation to convert intermediate representations into a distributions over tokens (see also Din et al. (2023)). By studying the layer-to-layer dynamics of this distribution, the authors noted that it tended to converge. This convergence is very suggestive that that the deeper layers could be prunable, while the fact that they had to train an affine probe is likely related to our observation that the final layer cannot be pruned. Somewhat relatedly, Gurnee and Tegmark (2023) observed that geographic features in the underlying text can be determined from linear probes trained on intermediate activations, as long as the activations are deeper than halfway.
More abstractly, Voita et al. (2023) and Liu et al. (2023b) found that the sparsity of activations transitions at around halfway through a network’s forward pass, evolving from sparse to dense. Perhaps relatedly, Panigrahi et al. (2023) investigated which model weights update the most during finetuning, finding that it’s those in the mid-layers.
Appendix B Experimental Details
Here we explain various details of models and healing (§B.1) and of evaluations (§B.2).
B.1 Model and healing details
@@
  • For quantization, we used the bitsandbytes library for QLoRA (Dettmers et al., 2023) to quantize our models to 4 bits.
@@
  • Also following Lee et al. (2023), we only applied LoRA to FFN modules: ["gate_proj", "down_proj", "up_proj"] for Llama-2 and Mistral models, ["fc1", "fc2"] for Phi-2, and ["w1", "w2", "c_proj"] for Qwen models.
The large majority of these hyperparameter choices are standard and found in previous works, e.g. Lee et al. (2023) and Dettmers et al. (2022). For absolute clarity, we list display all the model specific architecture and healing details below: